∫ [x 2 /(x 4 +1)]d[(x-1)/x]

∫ [x2 /(x4+1)]d[(x-1)/x]

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear upasana,

d[(x-1)/x]=1/x2dx
so the integral becomes integral of 1/1+x^4.

If we integrate this we get= 1/(4 Sqrt(2))(-2ArcTan(1-Sqrt(2)x) + 2ArcTan(1+Sqrt(2)x) - Ln| -1 +

Sqrt(2)x-x^2| + Ln| 1 + Sqrt(2)x+x^2|)

You can check from http://integrals.wolfram.com/index.jsp

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

 

Sagar Singh

B.Tech IIT Delhi

sagarsingh24.iitd@gmail.com